OpenStudy (anonymous):

The midpoint of (8, 6) and (-2, 12) is (3, 9)?

OpenStudy (anonymous):

True/False

geerky42 (geerky42):

Midpoint of \((x_1,y_1)\) and \((x_2,y_2)\) is \(\left(\dfrac{x_1+x_2}{2},\dfrac{y_1+y_2}{2}\right)\)

geerky42 (geerky42):

Here, you are given points \((8,6)\) and \((-2,12)\)

geerky42 (geerky42):

\(x_1 = 8\\~\\x_2 = -2\\~\\y_1 = 6\\~\\y_2 = 12\) So you have \(\left(\dfrac{8+\text-2}2,\dfrac{6+12}2\right)\quad\Longrightarrow\quad ?\)

OpenStudy (anonymous):

its false

OpenStudy (anonymous):

thank you

geerky42 (geerky42):

Explain why @slade

OpenStudy (anonymous):

just look on a graph

geerky42 (geerky42):

I think you should take a second look at graph.